OpenStudy (anonymous):

If the volume of a sphere is 36 cubic units, what is the radius? How would I write the equation to solve this

OpenStudy (montesanna99):

Ur equation would be 36= 4/3*3.14*r^3

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }\pi r ^{3}=36\] \[[\frac{ 3 }{ 4\pi }36]^{1/3}=r\]

OpenStudy (anonymous):

So what would my answer be? I'm confused

OpenStudy (anonymous):

use the equation I have setup and solve for r

OpenStudy (anonymous):

Is the radius 9pi?

OpenStudy (anonymous):

not quite a number to a power can also be thought of as the cube root of a number

OpenStudy (anonymous):

a number to the 1/3 specifically

OpenStudy (anonymous):

So what would the next step be?

OpenStudy (anonymous):

@VeritasVosLiberabit

OpenStudy (anonymous):

use this equation and this will be your final answer \[r=\sqrt[3]{\frac{ 3(36) }{ 4\pi }}\]

OpenStudy (anonymous):

9 units?

OpenStudy (anonymous):

@VeritasVosLiberabit

OpenStudy (anonymous):

@elementwielder would 9 units be my answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

one second...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no i got r= 2.04835...

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

your welcome